d. 35.0 ml of H2SO4 requires 20.0 ml of 0.100M NaOH. Find M for H2SO4.
1
Expert's answer
2020-06-15T14:03:22-0400

d)Equation for the reaction:

"2NaOH_{aq}+{H_2SO_4}_{aq}- {Na_2SO_4}{aq}+2H_2O_{l}"

"2" mol of "NaOH" reacts with "1" mol. of "H_2SO_4"

Mol "NaOH" in "20.00ml" of "0.100M" solution "=" "20mL\over1000mL" "* 0.100M"

"=2.0*10^{-3} mol"

This reacts with "{2.0*10^{-3}}\over2" "=" "1.0*10^{-3}" "mol"

"\therefore" "35mL" of "H_2SO_4" solution contains "1.0*10^{-3}" mol of "H_2SO_4"

"1.0L" "=1000mL" solution contains;

"1000mL\over 35mL" * "1.0*10^{-3}mol" "=0.0286 mol"

Molarity of "H_2SO_4" "=0.0286M"